If y varies directly as d find the constant of variation in the proportion 4/v=2/d

Answer by Theo(13342)

y = kd

this means y varies directly with d.

equation is 4/y = 2/d

cross multiply to get 2y = 4d

divide both sides by 2 to get y = 2d

constant of variation is 2.

what this says is that d is always 2 times the value of y.

to confirm, start with your original equation of 4/y = 2/d

choose a value of d at random.

i chose 556.

since y = 2d, then y must be equal to 1112

your equation of 4/y = 2/d becomes 4/1112 = 2/556

simplify both to get 1/278 = 1/278

you can choose any number for d, and you will find that y will have to be 2 times that for the equation to be true.

another example:

start with 4/y = 2/d

let d = 4

equation becomes 4/y = 2/4

multiply both sides by y and multiply both sides by 4 to get 16 = 2y

divide both sides by 2 to get y = 8

8 is w times the value of 4.

this will always happen when y = 2d.

direct variation form is y = kd.

k is the constant of variation which is equal to 2.